Filter Wheel Noise/Clunk

Monday, June 6, 2016 10:41 AM

Problem(s) Encountered:

About half way through the night we encountered a filter wheel problem.

We were switching filters and the MIS reported a failed move and wouldn''t

allow other attempts.  The failure was accompanied by a loud CLUNK,

which we later verified was from the filter wheel.

Eventually, we found that reinitializing the MIS would give us control of

the filter wheel again.  Most times, we were able to move to a new filter

after doing this, but not always.  Invariably, a subsequent move would

fail and we''d have to repeat.  The loud clunks continued, whether a

filter change worked or not.  The noise and the failures happened no

matter which direction or how many steps we were trying to go.

We sent the telescope to zenith and looked in through the door of the

filter wheel.  You can see what''s making the noise (metal bar over to the

right as you''re looking through the door).  One thing we noticed when

shining a flashlight in through the door is that the filters look FILTHY.

Can they be cleaned?


Filters were actually not excessively dirty. Still, they were pulled, brushed and blown off.

The noise heard was the sound of the thin flex-coupling on the detent roller arm assembly shearing (see image). It will be sent back to OSU, who is also making a spare (or two?) for MDM. In the meantime, the COM electronics was pulled from the 1.3m FW and installed on the 2.4m FW which was then brought down and installed on the 1.3m telescope. Camera and electronics were installed and all filters inserted. All systems tested ok.

Since the FW isn''t needed on the 2.4m, at least through shut down, we have some breathing room. Also incidentally, lost a helicoil on the 1.3m FW when removing it from the telescope.